# Hypergraph Neural Network with State Space Models for Node Classification

### v1

@@ 3.1.2 Node-Link Hypergraph
This approach creates a hypergraph $G^{l}_{h}$ by leveraging the neighboring connections within a graph $G$. In $G^{l}_{h}$, the neighboring nodes of a central node are clustered together to create a hyperedge. $H^{l}\in\mathbb{R}^{N\times N^{l}_{E}}$ represents the resulting hypergraph structure, where $N$ denotes the total number of nodes in $G$, and $N^{l}_{E}$ represents the number of hyperedges. The degree of nodes and hyperedges within $H^{l}$ are denoted by $D^{l}_{v}$ and $D^{l}_{e}$, respectively.
3.1.3 Hypergraph Based on Node Degree
According to their degrees, nodes are grouped to form a hypergraph $G^{d}_{h}$. Nodes sharing the same degree are combined into a single hyperedge. The resulting hypergraph structure is represented as $H^{d}\in\mathbb{R}^{N\times N^{d}_{E}}$, where $N$ is the total number of nodes, and $N^{d}_{E}$ corresponds to the number of unique node degrees. The degree matrices for nodes and hyperedges in $H^{d}$ are denoted as $D^{d}_{v}$ and $D^{d}_{e}$, respectively.
@@ 3.2 Mamba workflow
HGMN incorporates an attention mechanism to enable the adaptive integration of role-based and adjacency-based representations, enhancing performance in downstream tasks. Role features are embedded into $\hat{X}_{r}$ to achieve this. Additionally, HGMN leverages a structured state space model (SSM) attention mechanism for effective sequence modeling, ensuring the fusion of dynamic role and adjacency information.
SSM [18], a sequential model, employs a linear ordinary differential equation (ODE) to transform an input sequence $x(t)\in\mathbb{R}^{N}$ into an output sequence $y(t)\in\mathbb{R}^{N}$ through a latent state $h(t)\in\mathbb{R}^{N}$:
@@
where the state matrix is represented by $A\in\mathbb{R}^{N\times N}$, the input matrix, and the output matrix by $B$ and $C\in\mathbb{R}^{N}$. To determine $y(t)$ at time $t$, the latent state $h(t)$ must be solved, which is analytically challenging. Considering that real-world data is often discrete, we discretize the ODE as follows:
@@
where the discretized state parameters are $\bar{A}=\exp(\Delta\cdot A)$ and $\bar{B}=(\Delta\cdot A)^{-1}(\exp(\Delta\cdot A)-I)\cdot\Delta B$, with $\Delta$ standing for the step size.
In HGMN, the SSM attention mechanism utilizes these state transitions to integrate and propagate role and adjacency-based features effectively. By calculating the output sequence $y_{t}$, the model learns dynamic interactions between roles and structural features, enhancing its ability to represent higher-order relationships within the hypergraph. The attention coefficient $y_{r}$ is computed as follows:
@@
Eq (5) is used to generate the resulting fused embedding, $X_{f}\in\mathbb{R}^{N\times F_{h}}$, which is used as the input feature set for the hypergraph convolutional networks.
@@ 3.3 Hypergraph Convolution Layer
The fused features obtained from the attention mechanism (eq (5)) are the initial input of the hypergraph convolutional network. The node fusion features $X$, processed through a hypergraph convolutional network (HGMN), are computed as follows:
@@
Here:
  • 1.
    $W_{H}$ represents the weight matrix associated with the hyperedges.
  • 2.
    $D_{e}$ and $D_{v}$ represent the degree matrices for hyperedges and nodes, respectively.
  • 3.
    $H$ denotes the matrix representing the hypergraph structure.
  • 4.
    $X^{(l-1)}$ represents the input at the $l$-th layer, and $X^{(l)}$ corresponds to its output.
  • 5.
    $\delta^{(l-1)}$ denotes the learnable parameters for the $l$-th layer.
  • 6.
    $\sigma$ is the nonlinear activation function applied after the convolution operation.
@@ 3.4 Residual Network
A residual connection is incorporated to mitigate the over-smoothing problem often encountered in GNNs. This approach involves down-sampling the raw input features and combining them with the output from the last HGMN layer. The process is defined as:
@@
where $W_{\text{res}}\in\mathbb{R}^{F_{h}\times F_{h}}$ is the weight matrix for transformation, $X_{1}\in\mathbb{R}^{N\times F_{h}}$ represents the raw input features, $X_{\text{last}}\in\mathbb{R}^{N\times F_{h}}$ is the output from the last HGMN layer, $J\in\mathbb{R}^{N\times F_{h}}$ is the resulting representation after applying the residual connection.
@@ 3.5 Classification
The representation $J$ is passed through a fully connected layer followed by a softmax function to generate predictions $\hat{Y}$:
@@
where $b_{m}\in\mathbb{R}^{M}$ represents the bias term and $W_{m}\in\mathbb{R}^{F_{h}\times M}$ is the weight matrix.
The model’s loss function is defined as:
@@
where $\lambda$ is the regularization parameter, which helps prevent overfitting, $\delta$ denotes the model parameters.
@@ 4 Experiments
This section presents a series of comprehensive experiments across datasets of varying scales and methods to evaluate the effectiveness of the proposed HGMN model.
@@
In the ENZYMES dataset, nodes are categorized into primary and secondary types as outlined in [47]. The IIP dataset highlights extensive market connections between established corporations and emerging startups [48]. The TerroristRel dataset reflects disparities in the influence of different factions [49]. Meanwhile, in Cora, Citeseer, and Pubmed, key research papers serve as central references, being frequently cited by others [50].
Baselines: The performance of HGMN is benchmarked against the following well-established methods:
@@
    Role: Node role representations generated using GraphWave [11] are utilized as input features for classification with an SVM employing a Radial Basis Function (RBF) kernel [51].
@@
    StrucGCN [56]: StrucGCN model leverages a structural matrix based on topological similarity to enhance structural information learning, enabling effective non-local aggregation and improved performance on heterophily graphs.
@@
    AGNN [57]: AGNN integrates graph convolutional layers (GCL) with Graph Embedding Layers (GEL) derived from Laplacian-regularized optimization to mitigate over-smoothing by alternating between low- and high-order feature spaces.
Parameter Settings: In HGMN, we utilize two hypergraph convolutional layers followed by a fully connected layer. In Node2vec, the parameters $p$ and $q$ are both set to $1.0$. Adam optimizer is used with an initial learning rate of $0.003$, and the weight decay coefficient is set to 5e45e-4. For the baseline methods, the parameters are configured using their default settings.
Table 1: Performance of the proposed HGMN models in node classification, compared to the baseline models.

### v2

@@ 3.1.1 Node-Link Hypergraph
The first approach leverages the local neighborhood of each node in the input graph $G$. Specifically, the neighbors of a central node are grouped together to form a hyperedge, producing the hypergraph $G^{l}_{h}$. The resulting incidence matrix is represented as $H^{l}\in\mathbb{R}^{N\times N^{l}_{E}}$, where $N$ is the number of nodes in $G$ and $N^{l}_{E}$ the number of generated hyperedges. The associated degree matrices are denoted by $D^{l}_{v}$ and $D^{l}_{e}$. This construction emphasizes local connectivity by clustering nodes based on shared adjacency relations.
3.1.2 Degree-Based Hypergraph
The second approach groups nodes according to their degree values. Nodes with identical degrees form a common hyperedge, creating the hypergraph $G^{d}_{h}$. The structure is encoded by the incidence matrix $H^{d}\in\mathbb{R}^{N\times N^{d}_{E}}$, where $N^{d}_{E}$ corresponds to the number of distinct node degrees in the graph. The degree matrices of this hypergraph are denoted as $D^{d}_{v}$ and $D^{d}_{e}$. This method highlights global structural roles by linking nodes with similar connectivity levels.
Together, these two construction methods enable HGMN to capture both local adjacency-based relationships and global role-based similarities, providing a richer foundation for subsequent feature aggregation in the hypergraph convolutional network. Once hypergraphs capture both local and global structural information, we employ the Mamba state-space model to fuse role-based and adjacency-based features while efficiently modeling long-range dependencies.
@@ 3.2 Mamba Workflow
A central challenge in graph representation learning is the effective integration of local adjacency-based information and global role-based features. Standard attention mechanisms, while powerful, are computationally expensive for large graphs and may not fully capture long-range dependencies. To address this, HGMN incorporates the Mamba state space model (SSM) [22], which provides a linear-time alternative to traditional attention while maintaining strong capability in modeling long sequences. This makes Mamba well-suited for handling the dynamic interactions between structural and role-based embeddings. Fig. 3 illustrates the integration of adjacency-based and role-based features through the Mamba transformer mechanism, where structured state-space layers with learnable parameters adaptively fuse the two representations into a unified embedding.
Figure 3: Illustration of the Mamba transformer mechanism within HGMN. The block takes adjacency-based features ($\hat{X}_{a}$) and role-based features ($\hat{X}_{r}$) as inputs, processes them through structured state-space layers with learnable parameters $(A,B,C)$, and adaptively fuses the resulting outputs using attention coefficients $(\hat{y}_{a},\hat{y}_{r})$. The fused embedding $X_{f}$ captures both structural and role-aware information, making the mechanism well-suited for integrating higher-order graph dependencies.
Formally, an SSM models sequential dependencies through a latent state that evolves according to a linear ordinary differential equation (ODE). For an input sequence $x(t)\in\mathbb{R}^{N}$, the hidden state $h(t)$ and output sequence $y(t)$ are defined as:
@@
where $A\in\mathbb{R}^{N\times N}$ is the state matrix, and $B,C\in\mathbb{R}^{N}$ are input and output matrices. Since solving this ODE directly is intractable, it is discretized for practical use:
@@
where the parameters $\bar{A}$ and $\bar{B}$ are derived from $A,B$, and the step size $\Delta$.
Within HGMN, the SSM adaptively fuses role-based and adjacency-based embeddings. The SSM acts as a lightweight yet expressive attention mechanism, learning how information should be propagated across different roles and structural contexts in the hypergraph. Specifically, the fusion is realized through attention coefficients applied to role and adjacency embeddings:
@@
where $\hat{X}_{r}$ and $\hat{X}_{a}$ are the role-based and adjacency-based embeddings, and $\hat{y}_{r},\hat{y}_{a}$ are the learned attention weights from the SSM. The fused embedding $X_{f}\in\mathbb{R}^{N\times F_{h}}$ then serves as the input to the hypergraph convolutional layers. With the fused embeddings from Mamba, the hypergraph convolution layer aggregates higher-order features to encode both local adjacency and global role-based interactions.
@@ 3.3 Hypergraph Convolution Layer
The fused embedding $X_{f}$, which the Mamba-based attention mechanism produces (Eq. (5)), serves as the input to the hypergraph convolutional network. Unlike standard GNN pipelines that rely solely on adjacency information, this design ensures that both role-based dependencies and structural connectivity patterns are jointly encoded before propagation. The hypergraph convolution layer then performs higher-order feature aggregation as follows:
@@
where $H$ is the incidence matrix that encodes node–hyperedge relations, and $D_{v},D_{e}$ are the degree matrices of nodes and hyperedges used for normalization. The weight matrix $W_{H}$ captures how information flows across hyperedges, while the learnable parameter $\delta^{(l-1)}$ adaptively modulates contributions from neighboring nodes. Finally, the nonlinearity $\sigma$ introduces expressive power by enabling the network to learn complex interactions.
Integrating Mamba before the convolution step allows the model to capture long-range and role-aware dependencies that traditional hypergraph convolutions alone cannot efficiently represent. The convolution layer then grounds these enriched features in the hypergraph structure, ensuring that role–adjacency interactions are propagated across higher-order neighborhoods. This two-stage design, Mamba-based fusion followed by hypergraph convolution, enables HGMN to balance global context modeling with structural consistency, ultimately leading to more accurate and robust graph representation learning. To further stabilize training and prevent feature over-smoothing during propagation, we incorporate a residual connection that preserves both original and enriched embeddings.
@@ 3.4 Residual Network
The residual connection in HGMN mitigates the over-smoothing effect commonly observed in deep GNNs to further stabilize training and preserve feature diversity. Without such a mechanism, repeated propagation across hyperedges can cause node embeddings to become indistinguishable, limiting discriminative power. In our design, the residual pathway not only preserves the raw input features but also complements the Mamba-based fusion, ensuring that both the original role–adjacency information and the enriched higher-order representations are retained. The residual operation is defined as:
@@
where $X_{1}\in\mathbb{R}^{N\times F_{h}}$ denotes the down-sampled raw input features, $X_{\text{last}}\in\mathbb{R}^{N\times F_{h}}$ is the output from the final HGMN layer, and $W_{\text{res}}\in\mathbb{R}^{F_{h}\times F_{h}}$ is a learnable transformation matrix. The resulting representation $J\in\mathbb{R}^{N\times F_{h}}$ integrates both the untransformed feature space and the deeply propagated features, striking a balance between global sequence modeling (via Mamba) and local structural aggregation (via hypergraph convolution). This residual design ensures that Mamba-enhanced long-range dependencies are not diluted during convolutional propagation, while also avoiding oversmoothing by maintaining a direct signal from the raw input. Finally, the aggregated and stabilized node representations are fed into a classification layer to predict labels based on the fused structural and role-aware information.
@@ 3.5 Classification
The final representation $J$, which the model produces after integrating Mamba-based sequence modeling, hypergraph convolution, and residual preservation, is passed to a fully connected layer followed by a softmax activation to generate the predicted class probabilities $\hat{Y}$:
@@
where $W_{m}\in\mathbb{R}^{F_{h}\times M}$ is the learnable weight matrix and $b_{m}\in\mathbb{R}^{M}$ is the bias term. This layer acts as the final mapping from the Mamba-enhanced, role- and adjacency-aware embeddings into the label space of $M$ classes.
To optimize the model, we employ a cross-entropy loss with an additional regularization term:
@@
where $\delta$ denotes the trainable parameters and $\lambda$ controls the strength of regularization. The regularization term mitigates overfitting, which is especially important given that Mamba’s expressive sequence modeling can capture complex long-range dependencies that may otherwise lead to over-parameterization.
By leveraging Mamba at earlier stages, the classifier operates on features that already encode dynamic role interactions and adjacency-aware structural signals. This ensures that the final softmax layer does not merely perform shallow discrimination, but instead makes predictions based on deeply fused, higher-order hypergraph representations. Consequently, the classification step benefits directly from Mamba’s capacity to retain global contextual information while the residual path safeguards local feature integrity.
@@
4 Theoretical Analysis
In this section, we discuss the Complexity Bound and Convergence Behavior of the proposed HGMN model, providing theoretical insights into its computational efficiency and optimization properties.
Theorem 4.1 (Complexity Bound of HGMN).
Let $G=(V,E)$ be a graph with $n=|V|$ nodes and $m=|E|$ hyperedges. Consider an HGMN with $L$ layers, hidden dimension $d$ for hypergraph convolution layers, and a mamba block of hidden dimension $d_{m}$. Let $\bar{k}$ denote the average number of neighbors (or hyperedges) per node. Then, the computational complexity of HGMN per forward pass is bounded by:
$\mathcal{O}\Big(L\cdot(nd^{2}+md^{2}+n\bar{k}dd_{m}+nd_{m}^{2})+ndc\Big),$ (10)
where $c$ is the output dimension of the final fully connected layer.
Proof.
Let $\mathbf{X}^{(l)}\in\mathbb{R}^{n\times d}$ denote the node feature matrix at layer $l$, and $\mathbf{H}^{(l)}\in\mathbb{R}^{m\times d}$ denote the hyperedge feature matrix. The hypergraph convolution in HGMN is expressed as:
$\mathbf{X}^{(l+1)}=\sigma\Big(\mathbf{D}_{v}^{-1/2}\mathbf{H}\mathbf{W}_{e}\mathbf{D}_{e}^{-1}\mathbf{H}^{\top}\mathbf{D}_{v}^{-1/2}\mathbf{X}^{(l)}\mathbf{W}^{(l)}\Big),$ (11)
where $\mathbf{H}$ is the incidence matrix, $\mathbf{D}_{v}$ and $\mathbf{D}_{e}$ are the node and hyperedge degree matrices, $\mathbf{W}^{(l)}\in\mathbb{R}^{d\times d}$ and $\mathbf{W}_{e}\in\mathbb{R}^{d\times d}$ are trainable weights, and $\sigma$ is the activation function. The multiplication $\mathbf{X}^{(l)}\mathbf{W}^{(l)}$ has complexity $\mathcal{O}(nd^{2})$, and the hyperedge aggregation $\mathbf{H}\mathbf{W}_{e}\mathbf{D}_{e}^{-1}\mathbf{H}^{\top}\mathbf{X}^{(l)}$ has complexity $\mathcal{O}(md^{2})$.
The mamba block applies a neighborhood-aware transformation on each node:
$\mathbf{X}_{m}^{(l)}=\phi\Big(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{X}_{j}^{(l)}\mathbf{W}_{m}\Big),$ (12)
where $\mathcal{N}(i)$ is the set of neighbors for node $i$, $\alpha_{ij}$ are attention coefficients, $\mathbf{W}_{m}\in\mathbb{R}^{d\times d_{m}}$, and $\phi$ is an activation. The aggregation over $\bar{k}$ neighbors for all $n$ nodes gives complexity $\mathcal{O}(n\bar{k}dd_{m})$, and the internal linear transformation in the mamba block contributes $\mathcal{O}(nd_{m}^{2})$.
Summing the hypergraph convolution, mamba block, and the final fully connected layer $\mathbf{X}^{(L)}\mathbf{W}_{o}$ of dimension $d\times c$, the total per-layer cost is
$\mathcal{O}\big(nd^{2}+md^{2}+n\bar{k}dd_{m}+nd_{m}^{2}\big),$ (13)
and for $L$ layers, adding the final layer, the total complexity becomes:
$\mathcal{O}\Big(L\cdot(nd^{2}+md^{2}+n\bar{k}dd_{m}+nd_{m}^{2})+ndc\Big).$ (14)
This establishes the computational complexity bound for HGMN with the mamba block. $\square$
Theorem 4.2 (Convergence of HGMN).
Let the HGMN model be trained with the loss function
$L(\Theta)=\lambda|\delta|^{2}-\sum_{v\in V}\sum_{i=1}^{M}Y\ln\hat{Y}$, (15)
where $\Theta$ denotes all trainable parameters, including those in the hypergraph convolution layers, mamba block, and output layer, $V$ is the set of nodes, $M$ is the number of classes, $\hat{Y}$ is the predicted probability of node $v$ belonging to class $i$, and $\delta$ represents the set of learnable weights in HGMN. Assume that $L(\Theta)$ is differentiable and its gradient is Lipschitz continuous with constant $L_{g}>0$. If the parameters $\Theta$ are updated using gradient descent:
$\Theta_{t+1}=\Theta_{t}-\eta\nabla_{\Theta}L(\Theta_{t}),$ (16)
with a learning rate $0<\eta<\frac{2}{L_{g}}$, then the sequence $\{\Theta_{t}\}$ converges to a stationary point $\Theta^{*}$, i.e.,
$\lim_{t\to\infty}|\nabla_{\Theta}L(\Theta_{t})|=0.$ (17)
Proof.
Since the HGMN loss function $L(\Theta)=\lambda|\delta|^{2}-\sum_{v\in V}\sum_{i=1}^{M}Y\ln\hat{Y}$ is differentiable, its gradient $\nabla_{\Theta}L(\Theta)$ exists for all $\Theta$. The $L_{2}$-regularization term $\lambda\|\delta\|^{2}$ is convex and smooth, while the cross-entropy term $\sum_{v\in V}\sum_{i=1}^{M}Y\ln\hat{Y}$ is also smooth and Lipschitz continuous with respect to $\Theta$. Therefore, the full loss $L(\Theta)$ is $L_{g}$-smooth, i.e.,
$|\nabla_{\Theta}L(\Theta_{1})-\nabla_{\Theta}L(\Theta_{2})|\leq L_{g}|\Theta_{1}-\Theta_{2}|,\quad\forall\penalty 10000\ \Theta_{1},\Theta_{2}.$ (18)
Applying the standard descent lemma, we have:
$L(\Theta_{t+1})\leq L(\Theta_{t})-\eta|\nabla_{\Theta}L(\Theta_{t})|^{2}+\frac{L_{g}\eta^{2}}{2}|\nabla_{\Theta}L(\Theta_{t})|^{2}=L(\Theta_{t})-\eta\left(1-\frac{L_{g}\eta}{2}\right)|\nabla_{\Theta}L(\Theta_{t})|^{2}.$ (19)
For a learning rate $0<\eta<\frac{2}{L_{g}}$, the coefficient $\eta\left(1-\frac{L_{g}\eta}{2}\right)>0$, which ensures that $L(\Theta_{t})$ monotonically decreases. Since $L(\Theta)$ is lower bounded (cross-entropy is bounded below by 0 and $L_{2}$-norm is non-negative), the sequence $\{L(\Theta_{t})\}$ converges.
Summing over iterations from $t=0$ to T1T-1 gives:
$\sum_{t=0}^{T-1}|\nabla_{\Theta}L(\Theta_{t})|^{2}\leq\frac{L(\Theta_{0})-L^{*}}{\eta\left(1-\frac{L_{g}\eta}{2}\right)},$ (20)
where $L^{*}$ is the infimum of $L(\Theta)$. This implies that
$\lim_{t\to\infty}|\nabla_{\Theta}L(\Theta_{t})|^{2}=0,$ (21)
i.e., the gradient norm converges to zero, and $\Theta_{t}$ converges to a stationary point of $L(\Theta)$. ∎
The two theorems provided for the proposed HGMN model, complexity bound and convergence behavior, play a crucial role in understanding and justifying its theoretical foundations. The complexity bound theorem quantifies the computational cost of HGMN, demonstrating that despite incorporating sophisticated components such as hypergraph convolution and the mamba block, the model scales efficiently with the number of nodes, edges, and node features. This ensures that HGMN remains practical for large-scale graph datasets. The convergence theorem establishes that the training procedure is theoretically sound, guaranteeing that gradient-based optimization of the HGMN loss function converges to a stationary point under standard smoothness assumptions. Together, these theorems provide both computational assurance and optimization reliability, reinforcing that HGMN is not only empirically effective but also theoretically well-grounded.
4.1 Limitations and Future Directions of our proposed HGMN model
While the proposed HGMN demonstrates strong performance and robustness across multiple benchmarks, it is not without limitations. First, the current framework relies on pre-defined hypergraph construction strategies (node degree–based and neighborhood-level), which, although effective, may restrict adaptability when applied to datasets with substantially different structural properties. Future work can address this limitation by developing end-to-end hypergraph structure learning frameworks that allow the model to infer optimal hypergraph connectivity directly from data.
Second, our analysis is limited to static graphs. Many real-world systems, such as citation networks, communication systems, and biological interactions, evolve over time. Extending HGMN to handle dynamic or temporal hypergraphs would significantly broaden its applicability and improve its relevance to practical scenarios.
Finally, although we have provided both theoretical and empirical insights into computational complexity, scaling HGMN to extremely large graphs remains a challenge. Exploring approximation techniques, sampling strategies, or sparsity-inducing mechanisms may further improve scalability without sacrificing accuracy.
@@ 5 Experiments
This section presents a series of comprehensive experiments across datasets of varying scales and frameworks to evaluate the effectiveness of the proposed HGMN model.
@@
  7. 7.
    OGB-arxiv [58]: This dataset is a citation network derived from the Computer Science arXiv. Nodes correspond to arXiv papers, and directed edges denote citation relationships between them. Each paper is represented by a 128-dimensional word embedding of its title and abstract, and nodes are categorized into 40 subject areas.
  8. 8.
    OGB-products [58]: This dataset represents an Amazon product co-purchasing network. Nodes correspond to products, and edges denote that two products are frequently bought together. A 100-dimensional bag-of-words describes each product feature extracted from its reviews, and nodes are classified into 47 product categories.
Baselines: The performance of HGMN is benchmarked against the following well-established models:
@@
    Role: Node role representations generated using GraphWave [12] are utilized as input features for classification with a Support Vector Machine (SVM) employing a Radial Basis Function (RBF) kernel [59].
@@
    StrucGCN [64]: Structural Enhanced GCN (StrucGCN) model leverages a structural matrix based on topological similarity to enhance structural information learning, enabling effective non-local aggregation and improved performance on heterophily graphs.
@@
    AGNN [65]: Alternating Graph-Regularized Neural Network (AGNN) integrates graph convolutional layers (GCL) with Graph Embedding Layers (GEL) derived from Laplacian-regularized optimization to mitigate over-smoothing by alternating between low- and high-order feature spaces.
We further benchmark the proposed HGMN model against a range of state-of-the-art graph learning frameworks, including Do Transformers Really Perform Bad for Graph Representation? (Graphormer) [19], Graph Representation frOm self superVised mEssage passing tRansformer (GROVER) [20], Self-Supervised Masked Graph Autoencoders (GraphMAE) [21], Spectral Attention Network (SAN) [14], and Recipe for a General, Powerful, Scalable Graph Transformer (GPS) [18], on the large-scale OGB datasets.
Parameter Settings: In our experiments, datasets are divided into training, validation, and test splits using a 70/15/15 ratio with stratified sampling to preserve class balance. All results are reported as averages over five independent runs with different random seeds to ensure reproducibility and robustness. For the proposed HGMN model, we employ two hypergraph convolutional layers, followed by a fully connected layer to integrate both high-order structural dependencies and node-level features. A dropout rate of 0.5 is applied to prevent overfitting. Training is performed for up to 200 epochs using the Adam optimizer with a learning rate of 0.003 and a weight decay of $5\times 10^{-4}$. For Node2vec, the return parameter $p$ and in-out parameter $q$ are both set to 1.0, ensuring unbiased random walk exploration. Both $F_{r}$ and $F_{a}$ are fixed at 128, while the hidden size $F_{h}$ is selected from the range $\{60,80,100,120,140\}$. Hyperparameters for baseline frameworks are set according to their recommended configurations in the original implementations.
Table 2: Performance of the proposed HGMN models in node classification, compared to the baseline models. Values are reported as Micro-F1 $\pm$ std.